# Qianli Wang1,2   Mingyang Wang4,5,6   Nils Feldhus1,2,8   Simon Ostermann2,3,7 Yuan Cao9   Hinrich Schütze4,6   Sebastian Möller1,2   Vera Schmitt1,2 1Quality and Usability Lab, Technische Universität Berlin  2German Research Center for Artificial Intelligence (DFKI)

### v1

@@ Abstract
Quantization methods are widely used to accelerate inference and streamline the deployment of large language models (LLMs). While prior research has extensively investigated the degradation of various LLM capabilities due to quantization, its effects on model explainability and interpretability, which are crucial for understanding decision-making processes, remain unexplored. To address this gap, we conduct comprehensive experiments using three common quantization techniques at distinct bit widths, in conjunction with two explainability methods, counterfactual examples and natural language explanations, as well as two interpretability approaches, knowledge memorization analysis and latent multi-hop reasoning analysis. [^1] Our code will be publicly available once the paper is published. We complement our analysis with a thorough user study, evaluating selected explainability methods. Our findings reveal that, depending on the configuration, quantization can significantly impact model explainability and interpretability. Notably, the direction of this effect is not consistent, as it strongly depends on (1) the quantization method, (2) the explainability or interpretability approach, and (3) the evaluation protocol. In some settings, human evaluation shows that quantization degrades explainability, while in others, it even leads to improvements. Our work serves as a cautionary tale, demonstrating that quantization can unpredictably affect model transparency. This insight has important implications for deploying LLMs in applications where transparency is a critical requirement.
[Uncaptioned image] Through a Compressed Lens: Investigating the Impact of Quantization on LLM Explainability and Interpretability
@@
Figure 1: The effect of quantization on counterfactual examples (§5.1.1), natural language explanations (§5.1.1), knowledge memorization analysis (§5.1.2), and latent multi-hop reasoning (§5.1.2).
@@ 1 Introduction
With the shift towards LLMs and the focus on inference acceleration enabling more accessible and efficient deployment of LLMs, numerous quantization techniques have been developed Dettmers et al. (2022); Frantar et al. (2023); Xiao et al. (2023); Lin et al. (2024). By reducing the precision of a model’s parameters and lowering the number of bits used, quantization allows to decrease model size while mostly preserving its performance Gray and Neuhoff (1998). To date, the effects of quantization are evaluated across various aspects, including multilinguality Marchisio et al. (2024), bias Gonçalves and Strubell (2023), trustfulness Hong et al. (2024), alignment Jin et al. (2024), and instruction following capability Liu et al. (2024). Quantization has recently been ubiquitously employed for purposes of explainability and interpretability, aiming to explore black-box LLMs in areas such as parametric knowledge Namburi et al. (2023), natural language explanation Wang et al. (2024); Liu et al. (2024), and counterfactual example Bhattacharjee et al. (2024); Giorgi et al. (2025); Wang et al. (2025b). Despite the valuable insights that explainability and interpretability offer, the impact of quantization on explanation methods and explanation quality remains underexplored and has yet to be comprehensively characterized.
To bridge this gap, we present a comprehensive study that evaluates the effect of quantization on the performance and behavior of explainability and interpretability methods for LLMs. We first identify two explainability methods – counterfactual examples and natural language explanations – and two interpretability methods – knowledge memorization analysis and latent multi-hop reasoning analysis – and evaluate them on three widely used model quantization techniques (Figure 1). We conduct empirical experiments to evaluate four explainability and interpretability methods across their corresponding datasets, using four models of varying sizes in full precision and under three quantization techniques at different bit widths. We find that quantization leads to information loss within the model, and this trend becomes more pronounced in smaller models. Moreover, larger quantized models do not consistently outperform smaller full-precision models and quantization-induced degradation on explanation quality depends on the specific quantization techniques, and models deployed. Generally, integer quantization bib8 yields superior explanation quality compared to other quantization techniques across the evaluated methods. Surprisingly, lower-precision LLMs can sometimes surpass their higher-precision counterparts and a quantized model may even generate better explanations than the same model in full precision.
Secondly, we perform a user study to evaluate two selected explainability methods, assessing the explanations generated by models with different precisions for trustworthiness and coherence (Figure 1). Through the user study, we observe that explanations generated by full-precision models are perceived as slightly more trustworthy and coherent than those produced by quantized models. Meanwhile, automatic evaluation metrics for both explainability methods are not fully reliable and tend to overestimate the extent of explanation quality degradation resulting from quantization.
2 Preliminaries
@@
The decoding stage during LLM inference is typically memory-bound, where the key-value cache (KV cache) overhead often exceeds the size of the model weights Li et al. (2024c). Quantization techniques compress LLMs by converting model weights, activations, or the KV cache, originally in 32-bit floating-point format, into lower-precision data types Zhu et al. (2024), e.g., 8-bit integer Dettmers et al. (2022). These techniques can be broadly categorized into two types: quantization-aware training (QAT) and post-training quantization (PTQ). QAT requires retraining to mitigate errors introduced by quantization, whereas PTQ facilitates the direct application of a quantized model during inference. Thus, PTQ is especially appealing, as it offers computational efficiency without necessitating modifications to the model architecture or training process. Among PTQ, weight-only quantization is the most conventional and widely adopted method Wan et al. (2024); Zhou et al. (2024), which effectively accelerates matrix multiplications during the decoding stage Li et al. (2024c). Thereby, in this paper, we primarily evaluate the impact of weight-only quantization (§3.3) on explainability and interpretability (§4), which eliminates the need for retraining to address errors resulting from quantization.
Impact of Quantization.
Recent work has extensively examined the impact of quantization on various capabilities of LLMs. Marchisio et al. (2024) conduct a thorough analysis of quantized multilingual LLMs, focusing on performance degradation across languages. Gonçalves and Strubell (2023); Kirsten et al. (2024) explore the emergence of bias in the outputs generated by quantized models. Liu et al. (2024) find that in-context learning ability gradually declines in heavily quantized LLMs. Jin et al. (2024) observe that models with 4-bit quantization can still retain the alignment ability. In our work, we explicitly explore the impact of quantization on both interpretability and explainability.
Explainability & Interpretability.
Both explainability and interpretability focus on understanding how LLMs arrive at their decisions and how to make those decisions more understandable to humans. Although the distinction between these terms can be subtle and they are often used interchangeably Madsen et al. (2022); Lyu et al. (2024), the nuances between them are outlined as follows: Interpretability refers to the understandability of a model’s internal structure being inherently understandable, whereas explainability involves post-hoc techniques that analyze a model’s behavior to provide insights into its decisions Zhao et al. (2024); Calderon and Reichart (2025). Hendriks et al. (2025) experiment with several distillation methods, exploring their impact on model explainability in terms of natural language explanations. Their approach to studying the effect of explainability is most aligned with our work. In this study, we investigate how various quantization methods affect model explainability and interpretability.
@@ 3 Experimental Setup
We examine four representative methods (§3.1) to evaluate the impact of quantization on explainability and interpretability. Specifically, we compare full-precision LLMs (§3.4) with LLMs quantized using different techniques and varying bit configurations (§3.3) across multiple datasets (§3.2).
@@ 3.1 Methods
3.1.1 Explainability
Counterfactual Examples (CFE).
CFEs refers to edited inputs with minimal modifications that result in different model predictions Miller (2019); Madsen et al. (2022), which can be used for improving model performance Kaushik et al. (2020) and robustness Ross et al. (2021), and understanding the black-box nature of models Wu et al. (2021). To minimize the influence of factors such as in-context learning ability degradation Liu et al. (2024), we generate CFEs in a zero-shot setting using FIZLE Bhattacharjee et al. (2024).
Natural Language Explanations (NLE).
NLEs are free-text explanations that elucidate the prediction made by the model, which humans can easily understand Camburu et al. (2018); Wiegreffe et al. (2021); Wang et al. (2025a). Analogous to the CFEs, we generate NLEs in a zero-shot setting using ZeroCoT Kojima et al. (2022).
3.1.2 Interpretability
@@
As a first interpretability technique, we investigate the memorization of a model’s factual knowledge, identifying the reasons behind potential factual forgetting Namburi et al. (2023). By leveraging the theory of knowledge neurons, which suggests that specific neurons in LLMs are responsible for storing specific pieces of knowledge, we explore how quantization alters the storage and retrieval processes within these knowledge neurons. The findings aim to shed light on the trade-offs between model efficiency and fidelity to factual information, providing deeper insights into the impact of quantization on the internal mechanisms of LLMs.
@@
We reproduce the study on inspecting latent multi-hop reasoning errors by Yang et al. (2024). In summary, their analysis tests if LLMs use latent reasoning pathways and internal recall of knowledge about entities to answer complex factual queries. Since the authors reported a clear scaling trend for the first hop of reasoning but not for the second hop, we aim to investigate if this pattern can also be found for full-precision and quantized LLMs.
@@ 3.2 Datasets
Our study employs three widely recognized datasets [^2] Dataset examples and label distributions are detailed in Appendix A. commonly used for evaluating selected explainability and interpretability methods (§3.1).
eSNLI
Camburu et al. (2018) extends the SNLI corpus Bowman et al. (2015), which categorizes the relationship between a premise and a hypothesis into entailment, contradiction, or neutrality, by adding human-annotated NLEs.
LRE
Hernandez et al. (2024) is a knowledge probing dataset consisting of knowledge triplets $(s,r,o)$, structured in a one-hop setting, where $s$, $r$, and $o$ represents the subject, relation, and object, respectively.
TwoHop-Fact
Yang et al. (2024) is a dataset consisting of pairs of prompts: two-hop prompts ($\tau_{2H}$) for compositional queries, representing fact composition queries in the form $((e_{1},r_{1},e_{2}),(e_{2},r_{2},e_{3}))$, where $r_{1}$ and $r_{2}$ are relations and $e_{i}$ denotes an entity; and one-hop prompts ($\tau_{1H}$) for subqueries (Figure 7). $e_{1}$ and $e_{2}$ (in the second triplet) are subjects, while $e_{2}$ (in the first triplet) and $e_{3}$ objects. $e_{2}$ serves as a bridge entity, linking the two triplet to form a coherent two-hop reasoning chain.
@@ 3.3 Quantization Techniques
  • GPTQ Frantar et al. (2023) uses a second-order, Hessian-based optimization to quantize weights post-training with minimal accuracy loss;
  • AWQ Lin et al. (2024) enhances weight quantization by identifying and handling activation outliers to preserve model accuracy at low bit-widths;
  • Integer quantization Dettmers et al. (2022) implemented by BitsAndBytes [^3] The implementation of integer quantization provided in BitsAndBytes is limited to weight-only quantization: https://github.com/bitsandbytes-foundation/bitsandbytes. (bib4 and bib8) enables fast and memory-efficient inference by using optimized low-bit kernels.
@@ 3.4 Models
We deploy four instruction-tuned LLMs from Qwen2.5 family Qwen et al. (2024) with model sizes ranging from 7B to 72B (Table 1), across all explainability and interpretability methods (§3.1). These models are selected because their corresponding quantized versions are officially provided [^4] https://huggingface.co/collections/Qwen/qwen25-66e81a666513e518adb90d9e, ensuring the reproducibility of our results.
4 Evaluation
4.1 Automatic Evaluation
4.1.1 Explainability
Counterfactual Examples.
We evaluate CFEs using three widely used metrics: Label flip rate (LFR), perplexity (PPL), and text similarity (TS) Nguyen et al. (2024); Wang et al. (2025b). LFR captures the frequency with which the generated CFEs alter the model prediction to a different one. PPL assesses the fluency of the generated counterfactuals by measuring how well the model predicts the next word given the previous words. As CFEs should closely resemble the original inputs, TS is calculated using normalized word-level Levenshtein distances to quantify the similarity between the generated CFEs and the original inputs.
Natural Language Explanations.
NLEs are evaluated using three metrics: Micro-$F_{1}$ Score, BARTScore, and TIGERScore. Following the predict-then-explain pattern Kumar and Talukdar (2020), Micro-$F_{1}$ Score is used to assess task performance on the target dataset. BARTScore Yuan et al. (2021) is a reference-based metric that employs BART Lewis et al. (2020) to evaluate generated explanations based on how well they align with human-annotated references. TIGERScore Jiang et al. (2024), in contrast, is a reference-free metric that deploys a fine-tuned Llama2 Touvron et al. (2023) model to identify errors in the generated explanations and assigns a penalty score between $[-5,-0.5]$ for each mistake.
4.1.2 Interpretability
Knowledge Memorization Analysis
To assess the effects of quantization on factual knowledge, we first evaluate factual recall accuracy using the LRE dataset. A comparison of results before and after quantization reveals the degree of knowledge forgetting. We then employ the knowledge attribution method (Yu and Ananiadou, 2024) to trace the information loss back to specific layers and neurons, following the knowledge neuron theory. This combined analysis uncovers how quantization impacts the internal mechanisms responsible for storing and retrieving factual information.
Latent Multi-hop Reasoning Analysis (LMHR)
@@ 4.1.2 Interpretability
4.2 Human Evaluation
We further assess the effect of quantization on explainability methods and their quality (§3.1.1) by conducting a user study in which participants subjectively evaluate the explanations (CFEs and NLEs) based on two dimensions (§4.2.1). The goal is to investigate whether the explanation quality degradation observed through automatic evaluation aligns with the findings from the user study [^5] Human evaluation of interpretability methods is omitted, as internal model states, distributed knowledge representations, and latent reasoning paths are inherently abstract. These aspects lack clear ground truth for comparison and require specialized expertise to interpret and assess the explanations being provided Alangari et al. (2023); Zhu et al. (2025)..
4.2.1 Subjective Ratings
Following the design of Likert scales for explanation evaluation proposed by Feldhus et al. (2023); Chiang and Lee (2023), and user studies on CFEs and NLEs conducted by Domnich et al. (2025); Wang et al. (2025a), we ask human annotators to evaluate CFEs and NLEs based on the following dimensions, each rated on a 5-point Likert scale ranging from "strongly disagree" (1) to "strongly agree" (5):
  • Trustworthiness: Evaluate whether the provided explanation is truthfulness and can be relied upon by humans;
  • Coherence: Assess whether the provided explanation is sensible, clear, and coherent, and effectively captures the rationale.
Model NLE CFE
Preci Micro-$F_{1}$ BART TIGER LFR $\uparrow$ PPL $\downarrow$ TS $\downarrow$
-sion Score $\uparrow$ Score $\uparrow$ Score $\uparrow$
Qwen2.5-7B full 87.40% -6.56 -0.13 82.00% 94.34 0.35
bib4 84.80% -6.69 -0.22 75.00% 64.99 2.41
bib8 86.70% -6.56 -0.24 74.20% 88.85 1.61
gptq4 86.80% -6.53 -0.51 72.00% 92.39 0.36
gptq8 87.30% -6.60 -0.16 76.60% 79.48 1.63
awq 86.10% -6.59 -0.13 74.40% 99.97 0.40
Qwen2.5-14B full 69.50% -6.52 -0.25 87.20% 90.67 0.41
bib4 63.90% -6.73 -0.24 83.20% 93.03 0.44
bib8 67.70% -6.59 -0.21 85.00% 93.54 0.41
gptq4 78.50% -6.59 -0.25 86.80% 91.45 0.41
gptq8 67.90% -6.54 -0.30 85.40% 92.58 0.41
awq 70.00% -6.57 -0.81 86.60% 91.44 0.40
Qwen2.5-32B full 88.40% -7.68 -0.44 89.00% 87.07 0.40
bib4 89.50% -10.50 -1.28 87.80% 79.35 0.47
bib8 88.10% -8.77 -0.70 86.80% 86.11 0.40
gptq4 87.90% -6.61 -0.56 89.20% 87.07 0.40
gptq8 88.60% -7.91 -0.90 87.80% 90.03 0.39
awq 90.60% -9.91 -0.27 87.20% 92.82 0.38
Qwen2.5-72B full 78.80% -6.52 -0.47 89.60% 117.85 0.31
bib4 82.00% -6.54 -0.39 89.60% 120.36 0.33
bib8 78.90% -6.51 -0.52 88.40% 112.60 0.32
gptq4 78.00% -6.52 -0.47 87.80% 119.59 0.31
gptq8 78.10% -6.55 -0.53 90.00% 115.94 0.32
awq 77.70% -6.52 -0.51 88.20% 119.61 0.32
Table 1: Automatic evaluation results of NLEs and CFEs generated using ZeroCoT and FIZLE accordingly by Qwen2.5-{7B,14B,32B,72B} with full precision and different quantization methods on eSNLI, assessed by Micro-$F_{1}$ Score, BARTScore, and TIGERScore for NLEs; label flip rate (LFR), perplexity (PPL), and text similarity (TS) for CFEs. Bold-faced values indicate the best-performing approach for each respective model, while underlined values denote the best-performing quantization method.
4.2.2 User Study Setup
We conduct a user study involving $n=48$ participants, who are all native English speakers. We randomly sample ($k=30$) dataset indices. For each subset, i.e., model-precision pair, the explanations generated by the corresponding model, either in full precision or quantized using different methods (§3.3) with Qwen2.5-{7B,32B,72B}, are evaluated by at least two human annotators. We exclude Qwen2.5-14B due to its consistently suboptimal performance, both with and without quantization. Each annotator is assigned 15 explanations, accompanied by two evaluation dimensions (§4.2.1), and tasked with assigning appropriate scores based on a given Likert scale. We report inter-annotator agreements with Krippendorff’s $\alpha$ of $0.64$ for CFEs and $0.71$ for NLEs.

### v2

# Investigating The Impact of Quantization on Factual Knowledge Recall

@@ Abstract
Quantization methods are widely used to accelerate inference and streamline the deployment of large language models (LLMs). Although quantization’s effects on various LLM capabilities have been extensively studied, one critical area remains underexplored: factual knowledge recall (FKR), the process by which LLMs access stored knowledge. To this end, we conduct comprehensive experiments using three common quantization techniques at distinct bit widths, in conjunction with interpretability-driven analyses on two tasks, knowledge memorization and latent multi-hop reasoning.We show that quantization typically results in information loss within LLMs, consequently diminishing their capacity for FKR. This effect is particularly amplified in smaller models within the same architectural families. However, models quantized at reduced bit precision do not consistently exhibit inferior performance and occasionally quantization may even enhance model FKR. We find that BitSandBytes demonstrates highest preservation of the original full-precision model’s FKR. Despite variability across models and methods, quantization causes modest performance degradation and remains an effective compression strategy.
[Uncaptioned image] Through a Compressed Lens:
Investigating The Impact of Quantization on Factual Knowledge Recall
@@ 1 Introduction
The shift towards LLMs has created strong demand for efficient inference and accessible deployment. In response, numerous quantization techniques have been developed dettmers-etal-2022-8bits; frantar-etal-2023-optq; xiao-etal-2023-smooth; lin-etal-2024-awq. By reducing the precision of a model’s parameters, quantization allows us to decrease model size while mostly preserving its performance gray-etal-1998-quantization. Although the effects of quantization in LLMs have been evaluated across various aspects, e.g., multilinguality, bias, fairness, and trustworthiness marchisio-etal-2024-quantization; goncalves-strubell-2023-understanding; ramesh-etal-2023-comparative; hong-etal-2024-trust; liu-etal-2024-emergent, its impact on LLMs’ ability to recall factual knowledge from pretrained memory remains underexplored and has yet to be comprehensively characterized.
Figure 1: The effect of quantization on factual knowledge recall through knowledge memorization analysis and latent multi-hop reasoning analysis.
To this end, we present a comprehensive study of quantization’s impact on factual knowledge recall (FKR) in LLMs. Instead of solely observing the performance degradation, we perform interpretability-driven analyses on knowledge memorization and latent multi-hop reasoning3.1) tasks, to scrutinize behavior across neuron, layer, and model levels and evaluate three LLMs across three widely used model quantization techniques at different bit widths (Figure 1). We reveal that quantization typically impairs FKR, which is particularly acute for smaller models within a given model family. LLMs quantized to a lower bit precision do not consistently underperform those with higher precision. In some cases, quantization can even paradoxically enhance factual recall. Our findings indicate that BitSandBytes is most effective at preserving the model’s FKR. We observed no significant FKR degradation from quantization that would compromise model compression effectiveness, though effects vary by model and technique.
2 Background and Related Work
@@
Quantization techniques compress LLMs by converting model weights, activations, or the KV cache into lower-precision data types zhu-2024-surveymodelcompressionlarge. These techniques can be broadly categorized into two types: quantization-aware training (QAT) and post-training quantization (PTQ). QAT requires retraining to mitigate errors introduced by quantization, whereas PTQ facilitates the direct application of a quantized model during inference. In this paper, we primarily evaluate the impact of weight-only quantization (§3.3) on FKR (§4), which eliminates the need for retraining to address errors resulting from quantization.
Impact of Quantization. Recent work has extensively examined the impact of quantization on various capabilities of LLMs. marchisio-etal-2024-quantization conduct a thorough analysis of quantized multilingual LLMs, focusing on performance degradation across languages. goncalves-strubell-2023-understanding; kirsten-etal-2024-bias explore the emergence of bias in the outputs generated by quantized models. liu-etal-2024-emergent find that in-context learning ability gradually declines in heavily quantized LLMs. jin-etal-2024-comprehensive observe that models with 4-bit quantization can still retain the alignment ability. singh2025interpretingeffectsquantizationllms investigate the impact of quantization on model calibration. In our work, we explicitly explore how quantization affects FKR.
@@ 3 Experimental Setup
We examine two representative methods (§3.1) to evaluate the impact of quantization on FKR. Specifically, we compare full-precision LLMs (§3.4) with LLMs quantized using different techniques and bit configurations (§3.3) across two datasets (§3.2).
@@ 3.1 Methods
We investigate the memorization of a model’s factual knowledge, identifying the reasons behind potential factual forgetting namburi-etal-2023-cost. By leveraging the theory of knowledge neurons, which suggests that specific neurons in LLMs are responsible for storing specific pieces of knowledge, we explore how quantization alters the storage and retrieval processes within these neurons.
@@
We adapt the methodology from yang-2024-latent-multi-hop-reasoning for inspecting latent multi-hop reasoning errors. Specifically, the analysis tests whether full-precision and quantized LLMs employ similar latent reasoning pathways and internal knowledge recall mechanisms to answer complex factual queries.
@@ 3.2 Datasets
Our study employs two widely recognized datasets [^1] Dataset examples are detailed in Appendix A. for evaluating FKR with selected interpretability methods (§3.1).
LRE hernandez-2024-linearity is a knowledge probing dataset consisting of knowledge triplets $(s,r,o)$, structured in a one-hop setting, where $s$, $r$, and $o$ represents the subject, relation, and object, respectively.
TwoHop-Fact yang-2024-latent-multi-hop-reasoning is a dataset consisting of pairs of prompts: two-hop prompts ($\tau_{2H}$) for compositional queries, representing fact composition queries in the form $((e_{1},r_{1},e_{2}),(e_{2},r_{2},e_{3}))$, where $r_{1}$ and $r_{2}$ are relations and $e_{i}$ denotes an entity; and one-hop prompts ($\tau_{1H}$) for subqueries (Figure 4). $e_{1}$ and $e_{2}$ (in the second triplet) are subjects, while $e_{2}$ (in the first triplet) and $e_{3}$ objects. $e_{2}$ serves as a bridge entity, linking the two triplet to form a coherent two-hop reasoning chain.
@@ 3.3 Quantization Techniques
  • GPTQ frantar-etal-2023-optq uses a second-order, Hessian-based optimization to quantize weights post-training with minimal accuracy loss;
  • AWQ lin-etal-2024-awq enhances weight quantization by handling activation outliers to preserve model accuracy at low bit-widths;
  • Integer quantization dettmers-etal-2022-8bits implemented by BitsAndBytes (bib4 and bib8) enables fast and memory-efficient inference by using optimized low-bit kernels.
@@ 3.4 Models
We evaluate Llama3-8B llama3modelcard, Qwen2.5-7B and Qwen2.5-14B qwen2024qwen25technicalreport using two interpretability methods (§3.1). These models are selected due to the availability of quantized versions for each [^2] Detailed information for each model, including links to their respective quantized versions, is provided in Table 1., ensuring the reproducibility of our results.
(a) Top neuron distribution (Qwen2.5-7B)
(b) Top neuron distribution (Llama3-8B)
(c) Attention (Qwen2.5-7B): Landmark on continent
(d) FFN (Qwen2.5-7B): Landmark on continent
Figure 2: Top: Distribution of high-contributing neurons across layers, showing the average number of top-300 neurons per layer for Qwen2.5-7B (left) and Llama3-8B (right). Bottom: Layer-wise drop in neuron contribution scores across quantization methods for the landmark on continent relation in Qwen2.5-7B, comparing attention sublayers (left) and feed-forward sublayers (right).
4 Evaluation Setup
Knowledge Memorization Analysis.
To assess the effects of quantization on FKR, we first evaluate factual recall accuracy using LRE. A comparison of results before and after quantization reveals the degree of knowledge forgetting. We then employ a knowledge attribution method (yu-ananiadou-2024-neuron) to trace the information loss back to specific layers and neurons, following the knowledge neuron theory. This combined analysis uncovers how quantization impacts the internal mechanisms responsible for storing and retrieving information.
Latent Multi-hop Reasoning Analysis (LMHR).